OpenStudy (babyjayy):

Solve: negative 1 over 2 x + 1 = –x + 8 7 1 over 2 14 18

Mehek (mehek14):

Does it look like this: \(\bf \dfrac{1}{2}x+1=-x+8\) ?

OpenStudy (babyjayy):

the fraction is a negative

Mehek (mehek14):

But everything else is right?

OpenStudy (babyjayy):

yea

Mehek (mehek14):

Ok, so like this \(\bf - \dfrac{1}{2}x+1=-x+8\) You would move x to one side and the other number on the other side. Add x on both sides.

Mehek (mehek14):

Nope

OpenStudy (babyjayy):

dose x=14?

Mehek (mehek14):

Yes

OpenStudy (babyjayy):

can you help me with another one if i need help?

Mehek (mehek14):

Sure, but I have to go right now. :c

alones (alones):

@babyjayy make a new post ;) and i'll try assisting mehek